# Reinforced In-Context Black-Box Optimization

### v2

@@ 3.3 Data Generation
Data Collection. Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can select a diverse set of algorithms for data collection. For example, we can select some representatives from different types of BBO algorithms, e.g., BO and EA. Datasets $\mathcal{D}_{i,j}$ are obtained by using each behavior algorithm to optimize each function with different random seeds. Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented with RTG tokens $R_{t}$ computed as in Eq. (3). The resulting histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for model training.
Data Normalization. To provide a unified interface and balance the statistic scales across tasks, it is important to apply normalization to the inputs to our model. We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively. For the function value $y$, we apply random scaling akin to previous works [14, 64]. That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$. The RTG tokens are calculated accordingly with the normalized values. The random normalization can make a model exhibit invariance across various scales of $y$.
@@ 4 Experiments
In this section, we examine the performance of RIBBO on a wide range of tasks, including synthetic functions, Hyper-Parameter Optimization (HPO) and robot control problems. The model architecture and hyper-parameters are maintained consistently across these problems. We train our model using five distinct random seeds, ranging from 00 to $4$, and each trained model is run five times independently during the execution phase. We will report the average performance and standard deviation. Details of the model hyper-parameters are given in Appendix A. The codes are provided in the supplementary.
@@ 4.1 Experimental Setup
Benchmarks. We use BBO Benchmarks BBOB [15], HPO-B [2], and rover trajectory planning tasks [60]. The BBOB suite, a comprehensive and widely used benchmark in the continuous domain, consists of $24$ synthetic functions. For each function, a series of transformations are implemented on the search space to obtain a distribution of functions with similar properties. According to the properties of the functions, they can be divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. HPO-B is the most commonly used HPO benchmark and consists of a series of HPO problems. It fits an XGBoost model as the objective function for each HPO problem, and we conduct experiments on two widely used machine models, SVM and XGBoost. For robot control optimization, we perform experiments on rover trajectory planning, which is a trajectory optimization task to emulate rover navigation. Similar to [15, 59], we implement random translations and scalings to the search space to construct a distribution of functions. For BBOB and rover problems, we sample a set of functions as training and test tasks, while for HPO-B, we use the meta-training/test splits provided by the authors. Detailed explanations of the benchmarks can be found in Appendix B.1.
Data. Similar to OptFormer [14], we employ $7$ behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [50], Eagle Strategy [66], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively. Datasets are generated by using each behavior algorithm to optimize each training function with different random seeds. More details about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
@@ 4.2 Baselines
As RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with offline datasets, including Behavior Cloning (BC) [7] and OptFormer [14]. Their hyper-parameters are set as same as that of our model for fairness. Note that the seven behavior algorithms used to generate datasets are also important baselines, and included for comparison as well.
@@
OptFormer employs a transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution. Its original implementation is built upon a text-based transformer with a large training scale. In this paper, we re-implement a simplified version of OptFormer where we only retain the algorithm type within the metadata. Various initial states, denoted as $\bm{x}_{0}$ and $y_{0}$, are used to distinguish between algorithms. They are obtained by indexing the algorithm type through an embedding layer, thereby enabling the identification of distinct behavior algorithms within OptFormer. Further training details can be found in Appendix C.
@@ 4.3 Main Results
The results are shown in Figure 2. For the sake of clarity in visualization, we have omitted the inclusion of Random Search and Shuffled Grid Search due to their poor performance from start to finish. We can observe that RIBBO achieves superior or at least equivalent efficacy in comparison to the best behavior algorithm on each problem except SVM and rover. This demonstrates the versatility of RIBBO, while the most effective behavior algorithm depends upon the specific problem at hand, e.g., the best behavior algorithms on Lunacek, Rastrigin and XGBoost are GP-EI, Eagle Strategy and CMA-ES, respectively. Note that the good performance of RIBBO does not owe to the memorization of optimal solutions, as the search space is transformed randomly, resulting in variations in optimal solutions across different functions from the same distribution. It is because RIBBO is capable of using RTG tokens to identify algorithms and reinforce the performance on top of the behavior algorithms, which will be clearly shown later. We can also observe that RIBBO performs extremely well in the early stage, which draws the advantage from the HRR strategy, i.e., employing zero as the immediate RTG to generate the optimal potential solutions.
@@
Compared with BC and BC Filter, RIBBO performs consistently better except on the SVM problem. BC tends to imitate the average behavior of various algorithms, and its poor performance is due to the aggregation of data from behavior algorithms with inferior performance. BC Filter is generally better than BC, because the data from the two underperforming behavior algorithms, i.e., Random Search and Shuffled Grid Search, are manually excluded for the training of BC Filter. As introduced before, OptFormer requires to manually specify which algorithm to execute. We have specified the behavior algorithm Eagle Strategy in Figure 2, which obtains good overall performance on these problems. It can be observed that OptFormer displays a close performance to Eagle Strategy, while RIBBO performs better. More results about the imitation capacity of OptFormer can be found in Appendix C.
@@
Cross-Distribution Generalization. We further conduct experiments to examine the generalization of RIBBO to unseen function distributions. For this purpose, we train the model on $4$ of the $5$ chosen synthetic functions and test it with the remaining one. Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and we sample a set of functions from each distribution as introduced before. Greiwank Rosenbrock is used for testing. The results in Figure 3(b) suggest that RIBBO has a strong generalizing ability to unseen function distributions. More results on cross-distribution generalization are deferred to Appendix D. The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in [13]. In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
@@
Figure 3: (a) Visualization of the contour lines of $2$D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO. (b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties. The legend shares with that of Figure 2. (c) Initial RTG $R_{0}$’s influence on performance. (d) RTG update strategy comparison between HRR and the naive strategy with various $R_{0}$.
@@ 4.4 Ablation Studies
Influence of Initial RTG Token $R_{0}$. By incorporating RTG tokens, RIBBO can attend to RTGs and generate optimization trajectories based on the specified initial RTG token $R_{0}$. To validate this, we examine the performance of RIBBO with different values of $R_{0}$, and the results are presented in Figure 3(c). Here, the RTG values, instead of normalized objective values, are used as the $y$-axis. We can observe that the cumulative regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
Effectiveness of HRR. A key point of the inference procedure is how to update the value of RTG tokens at each iteration. To assess the effectiveness of the proposed strategy HRR in Eq. (5), we compare it with the naive strategy, that sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration. The results are shown in Figure 3(d). The naive strategy displays distinct behaviors when $R_{0}=0$ or $R_{0}=5$. When $R_{0}=0$, i.e., the lower bound of regret, the model performs well initially. However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values. Using $R_{0}=5$ compromises the initial performance, as the model may not select the most aggressive solutions with a high $R_{0}$. However, a higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in later stage. The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG to 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq. (5) ensures the values remain meaningful and feasible.
Further Studies. We also study the effects of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for $y$, the model size, and the sampled subsequence length $\tau$. Please see Appendix E. Generalization studies across the entire BBOB suite are shown in Appendix D, while the effects of transformations applied to the search space are detailed in Appendix F.
@@ 5 Conclusion
This paper proposes RIBBO, which employs a transformer architecture to learn a reinforced BBO algorithm from offline datasets in an E2E fashion. By incorporating RTG tokens into the optimization histories, RIBBO can automatically generate optimization trajectories satisfying the user-desired regret. Comprehensive experiments on BBOB, HPO and robot control problems show the versatility of RIBBO. This work is a preliminary attempt towards universal BBO, and we hope it can encourage more explorations in this direction. For example, we only consider the model training over continuous search space with the same dimensionality, and it would be interesting to explore heteroscedastic search space with different types of variables.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
@@ Appendix A Model Details
We employ the commonly used GPT architecture [48] and the hyper-parameters are maintained consistently across the problems. Details can be found in Table 1. For BC, BC Filter and OptFormer, the hyper-parameters are set as same as those of our model. The training takes about $20$ hours on $1$ GPU (Nvidia RTX $4090$).
@@ B.1 Benchmarks
  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in the continuous domain. This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties. According to the properties of functions, these synthetic functions can be divided into $5$ categories, i.e., (1) separable functions, (2) moderately conditioned functions, (3) ill-conditioned and unimodal functions, (4) multi-modal functions with adequate global structure, and (5) multi-modal functions with weak global structures. We select one function from each category to evaluate our algorithm, Rastrigin from (1), Rosenbrock Rotated from (2), Sharp Ridge from (3), Greiwank Rosenbrock from (4), and Lunacek from (5). We use the BBOB benchmark implementation in Open Source Vizier [^1] https://github.com/google/vizier, and the dimension is set to $10$ for all functions.
@@
  • SVM and XGBoost. HPO-B [2] is the most commonly used HPO benchmark and is grouped by search space id. Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost. Each such search space has multiple associated dataset id, which is a particular HPO problem, i.e., optimizing the performance of the corresponding model on a dataset. For the continuous domain, it fits an XGBoost model as the objective function for each HPO problem. These datasets for each search space id are divided into training and test datasets. We examine our method on two selected search space id, i.e., $5527$ and $6767$, which are two representative HPO problems tuning SVM and XGBoost, respectively. SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most in HPO-B. We use the official implementations [^2] https://github.com/releaunifreiburg/HPO-B.
@@
  • Rover Trajectory Planning [16, 60] is a trajectory optimization task designed to emulate a rover navigation task. The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, resulting in a total of $60$ parameters to optimize. Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{s}\|_{1}+\|\bm{x}_{58,59}-\bm{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of the given trajectory, $\bm{x}_{0,1}$ denotes the first and second dimensions of $\bm{x}$ (similarly for $\bm{x}_{58,59}$), $\bm{s}$ and $\bm{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem. This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions. To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ and scalings in $[0.9,1.1]$, similar to previous works [15, 59]. The training and test tasks are randomly sampled from the distribution. We use the standard implementation for the rover problem [^3] https://github.com/uber-research/TuRBO.

### v3

@@ 3.3 Data Generation
Data Collection. Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can employ a diverse set of behavior algorithms for data collection. For example, we can select some representatives from different types of BBO algorithms, e.g., BO and EA. Datasets $\mathcal{D}_{i,j}$ are obtained by using each behavior algorithm to optimize each task with different random seeds. Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented with RTG tokens $R_{t}$, which is computed as in Eq. (3). The resulting histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for model training.
Data Normalization. To provide a unified interface and balance the statistic scales across tasks, it is important to apply normalization to the inputs to our model. We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively. For the function value $y$, we apply random scaling akin to previous works Wistuba & Grabocka (2021); Chen et al. (2022). That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$ for training. The RTG tokens are calculated accordingly with the normalized values. The random normalization can make a model exhibit invariance across various scales of $y$. For inference, the average values of the best-observed and worst-observed values across the training tasks are used to normalize $y$.
@@ 4 Experiments
In this section, we examine the performance of RIBBO on a wide range of tasks, including synthetic functions, Hyper-Parameter Optimization (HPO) and robot control problems. The model architecture and hyper-parameters are maintained consistently across these problems. We train our model using five distinct random seeds, ranging from 00 to $4$, and each trained model is run five times independently during the execution phase. We will report the average performance and standard deviation. Details of the model hyper-parameters are given in Appendix A. Our code is available at https://github.com/songlei00/RIBBO.
@@ 4.1 Experimental Setup
Benchmarks. We use BBO Benchmarks BBOB Elhara et al. (2019), HPO-B Arango et al. (2021), and rover trajectory planning task Wang et al. (2018). The BBOB suite, a comprehensive and widely used benchmark in the continuous domain, consists of $24$ synthetic functions. For each function, a series of linear and non-linear transformations are implemented on the search space to obtain a distribution of functions with similar properties. According to the properties of these functions, they can be divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. HPO-B is a commonly used HPO benchmark and consists of a series of HPO problems. Each problem is to optimize a machine learning model across various datasets, and an XGBoost model is provided as the objective function for evaluation in a continuous space. We conduct experiments on two widely used models, SVM and XGBoost, in the continue domain. For robot control optimization, we perform experiments on rover trajectory planning task, which is a trajectory optimization problem to emulate rover navigation. Similar to Elhara et al. (2019); Volpp et al. (2020), we implement random translations and scalings to the search space to construct a distribution of functions. For BBOB and rover problems, we sample a set of functions from the task distribution as training and test tasks, while for HPO-B, we use the meta-training/test task splits provided by the authors. Detailed explanations of the benchmarks can be found in Appendix B.1.
Data. Similar to OptFormer Chen et al. (2022), we employ $7$ behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution Real et al. (2019), Eagle Strategy Yang & Deb (2010), CMA-ES Hansen (2016), and GP-EI Balandat et al. (2020), which are representatives of heuristic search, EA, and BO, respectively. Datasets are generated by employing each behavior algorithm to optimize various training functions sampled from the task distribution, using different random seeds. For inference, new test functions are sampled from the same task distribution to serve as the test set. Specifically, for the HPO-B problem, the meta-training/test splits have been predefined by the authors and we adhere to this standard setup. Additional information about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
@@ 4.2 Baselines
As RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with offline datasets, including Behavior Cloning (BC) Bain & Sammut (1995) and OptFormer Chen et al. (2022). Their hyper-parameters are set as same as that of our model for fairness. Note that the seven behavior algorithms used to generate datasets are also important baselines, and included for comparison as well.
@@
OptFormer employs a transformer to imitate the behaviors of a set of algorithms and an algorithm identifier usually needs to be specified manually during inference for superior performance. Its original implementation is built upon a text-based transformer with a large training scale. In this paper, we re-implement a simplified version of OptFormer where we only retain the algorithm identifier within the metadata. The initial states, denoted as $\bm{x}_{0}$ and $y_{0}$, are used to distinguish between algorithms. They are obtained by indexing the algorithm type through an embedding layer, thereby aligning the initial states with the specific imitated algorithm. This enables the identification of distinct behavior algorithms within the simplified OptFormer. Further details about the re-implementation can be found in Appendix C.
@@ 4.3 Main Results
The results are shown in Figure 2. For the sake of clarity in visualization, we have omitted the inclusion of Random Search and Shuffled Grid Search due to their poor performance from start to finish. We can observe that RIBBO achieves superior or at least equivalent efficacy in comparison to the best behavior algorithm on each problem except SVM and rover. This demonstrates the versatility of RIBBO, while the most effective behavior algorithm depends upon the specific problem at hand, e.g., the best behavior algorithms on Lunacek, Rastrigin and XGBoost are GP-EI, Eagle Strategy and CMA-ES, respectively. Note that the good performance of RIBBO does not owe to the memorization of optimal solutions, as the search space is transformed randomly, resulting in variations in optimal solutions across different functions from the same distribution. It is because RIBBO is capable of using RTG tokens to identify algorithms and reinforce the performance on top of the behavior algorithms, which will be clearly shown later. We can also observe that RIBBO performs extremely well in the early stage, which draws the advantage from the HRR strategy, i.e., employing 00 as the immediate RTG to generate the optimal potential solutions.
@@
Compared with BC and BC Filter, RIBBO performs consistently better except on the SVM problem. BC tends to imitate the average behavior of various algorithms, and its poor performance is due to the aggregation of behavior algorithms with inferior performance. BC Filter is generally better than BC, because the data from the two underperforming behavior algorithms, i.e., Random Search and Shuffled Grid Search, are manually excluded for the training of BC Filter. As introduced before, OptFormer requires to manually specify which algorithm to execute. We have specified the behavior algorithm Eagle Strategy in Figure 2, which obtains good overall performance on these problems. It can be observed that OptFormer displays a close performance to Eagle Strategy, while RIBBO performs better. More results about the imitation capacity of OptFormer can be found in Appendix C.
@@
Generalization. We also conduct experiments to examine the generalization capabilities of RIBBO. For this purpose, we train the model on all $24$ BBOB synthetic functions simultaneously with the results shown in Figure 3(b). To aggregate results across functions with different output scaling, we normalize all the functions adhering to previous literature Turner et al. (2020); Arango et al. (2021); Chen et al. (2022). The results suggest that RIBBO demonstrates strong generalizing to a variety of functions with different properties. Further experiments are conducted to examine the cross-distribution generalization to unseen function distributions. The model is trained on $4$ of the $5$ chosen synthetic functions and tested with the remaining one. Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and a set of functions is sampled from each distribution as introduced before. The results suggest that RIBBO has a strong generalizing ability to unseen function distributions. Due to the space limitation, the results and more details on cross-distribution generalization are deferred to Appendix D. The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in Chen et al. (2017). In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
@@
Figure 3: (a) Visualization of the contour lines of $2$D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO. (b) Generalization by training the model across all $24$ BBOB synthetic functions simultaneously. The results across functions with different output scaling are normalized to obtain the aggregate results. The legend shares with that of Figure 2. (c) Initial RTG $R_{0}$’s influence on performance. (d) RTG update strategy comparison between HRR and the naive strategy with various initial RTG $R_{0}$.
@@ 4.4 Ablation Studies
Influence of Initial RTG Token $R_{0}$. By incorporating RTG tokens, RIBBO is able to attend to RTGs and generate optimization trajectories based on the specified initial RTG token $R_{0}$. To validate this, we examine the performance of RIBBO with different values of $R_{0}$, and the results are presented in Figure 3(c). Here, the RTG values, instead of normalized objective values, are used as the $y$-axis. We can observe that the cumulative regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation. We also conduct experiments to explore the effect of varying the immediate RTG $R_{t}$, and it is observed that setting to a value larger than 00 will decrease the performance and converge to a worse value. Please see Appendix E.
Effectiveness of HRR. A key point of the inference procedure is how to update the value of RTG tokens at each iteration. To assess the effectiveness of the proposed strategy HRR outlined in Eq. (5), we compare it with the naive strategy, that sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration. The results are shown in Figure 3(d). The naive strategy displays distinct behaviors depending on the initial setting of $R_{0}$. Specifically, when $R_{0}=0$, i.e., the lower bound of regret, the model performs well initially. However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values. Using $R_{0}=5$ compromises the initial performance, as the model may not select the most aggressive solutions with a high $R_{0}$. However, a higher initial $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in later stage. The proposed HRR strategy consistently outperforms across the whole optimization stage, because setting the immediate RTG to 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens, as specified in Eq. (5), ensures that these values remain meaningful and feasible.
Further Studies. We also study the effects of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for $y$, the model size, and the sampled subsequence length $\tau$. Please see Appendix F. More visualizations illustrating the effects of random transformations on the search space are detailed in Appendix G.
@@ 5 Conclusion
This paper proposes RIBBO, which employs a transformer architecture to learn a reinforced BBO algorithm from offline datasets in an E2E fashion. By incorporating RTG tokens into the optimization histories, RIBBO can automatically generate optimization trajectories satisfying the user-desired regret. Comprehensive experiments on BBOB synthetic functions, HPO and robot control problems show the versatility of RIBBO. This work is a preliminary attempt towards universal BBO, and we hope it can encourage more explorations in this direction. For example, we only consider the model training over continuous search space with the same dimensionality, and it would be interesting to explore heteroscedastic search space with different types of variables. A mathematical theoretical analysis of the in-context learning capabilities based on RTG tokens is also of interest.
@@ Appendix A Model Details
We employ the commonly used GPT architecture Radford et al. (2018) and the hyper-parameters are maintained consistently across the problems. Details can be found in Table 1. For BC, BC Filter and OptFormer, the hyper-parameters are set as same as those of our model. The training takes about $20$ hours on $1$ GPU (Nvidia RTX $4090$).
@@ B.1 Benchmarks
  • BBOB Elhara et al. (2019) is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in the continuous domain. This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties. According to the properties of functions, these synthetic functions can be divided into $5$ categories, i.e., (1) separable functions, (2) moderately conditioned functions, (3) ill-conditioned and unimodal functions, (4) multi-modal functions with adequate global structure, and (5) multi-modal functions with weak global structures. We select one function from each category to evaluate our algorithm, Rastrigin from (1), Rosenbrock Rotated from (2), Sharp Ridge from (3), Greiwank Rosenbrock from (4), and Lunacek from (5). We use the BBOB benchmark implementation in Open Source Vizier [^1] https://github.com/google/vizier, and the dimension is set to $10$ for all functions.
@@
  • SVM and XGBoost. HPO-B Arango et al. (2021) is the most commonly used HPO benchmark and is grouped by search space id. Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost. Each such search space has multiple associated dataset id, which is a particular HPO problem, i.e., optimizing the performance of the corresponding model on a dataset. For the continuous domain, it fits an XGBoost model as the objective function for each HPO problem. These datasets for each search space id are divided into training and test datasets. We examine our method on two selected search space id, i.e., $5527$ and $6767$, which are two representative HPO problems tuning SVM and XGBoost, respectively. SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most in HPO-B. We use the official implementations [^2] https://github.com/releaunifreiburg/HPO-B.
@@
  • Rover Trajectory Planning Eriksson et al. (2019); Wang et al. (2018) is a trajectory optimization task designed to emulate a rover navigation task. The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, resulting in a total of $60$ parameters to optimize. Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{s}\|_{1}+\|\bm{x}_{58,59}-\bm{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of the given trajectory, $\bm{x}_{0,1}$ denotes the first and second dimensions of $\bm{x}$ (similarly for $\bm{x}_{58,59}$), $\bm{s}$ and $\bm{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem. This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions. To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ and scalings in $[0.9,1.1]$, similar to previous works Elhara et al. (2019); Volpp et al. (2020). The training and test tasks are randomly sampled from the distribution. We use the standard implementation for the rover problem [^3] https://github.com/uber-research/TuRBO.